# Leveraging Graph Neural Networks and Multi-Agent Reinforcement Learning for Inventory Control in Supply Chains

### v1

@@ 4.2 Sensitivity to Gaussian Noise
The standard deviation values were changed from 0.0, 0.1, 0.2, 0.5, 1.0 and 2.0. Figure 13(a) shows the cumulative profit of the agents at execution of the trained policies. The performance increases up to a standard deviation of 0.5. After which, when the noise intensity increases, the overall performance decreases. This occurs as when the standard deviation of the Gaussian distribution is small, there’s insufficient noise, leading to a lack of regularization. This can result in overfitting, limited exploration, and a tendency to settle into local optima. This is particularly detrimental in scenarios with large state spaces, where effective exploration becomes crucial for agents to learn optimal policies. Conversely, when the standard deviation is 1.0 and 2.0 the overall performance becomes worse as seen in Figure 13(a). This occurs when the noise intensity is too high, excessive noise can introduce significant randomness, increasing the variance in the value function which disrupts the learning process which may hinder convergence.
@@
(a) Execution curves at different levels of noise
(b) Change in entropy from 3 different noise levels
@@ 4.3 Scalability
Moreover, Figure 15, compares the mean training time per iteration in seconds both in terms of methodologies and number of agents. As expected, as the number of agents increases, the training time per iteration increases due to the increased complexity of the learning process. The methodologies leveraging the graph structure also have a higher mean training time per iteration because they require complex message passing between agents in the graph. G-MAPPO, P-GCN-MAPPO, Noisy P-GCN-MAPPO utilize a GNN which requires training alongside the policies. Therefore, extracting and processing information from the graph structure itself adds computational overhead compared to simpler MARL approaches. As the number of agents increases, the training time per iteration increases at a faster rate for the methodologies using GNNs as the communication and information processing becomes increasingly complex as the number of agents and complexity of the structure increases. However, Figure 15 shows that methods with an integrated pooling mechanism have faster training times than those without one. This further emphasises the scalability advantage of integrating a pooling mechanism compared to a standard GNN architecture in a MARL framework. The integration of the pooling mechanism reduces the dimensionality of the input to the central critic, reducing the computational overhead which leads to faster training times. Therefore, as opposed to naïvely concatenating information, MARL frameworks that employ smart information aggregation techniques can offer several advantages: reduced computational complexity, improved scalability and potential for improved performance by efficient aggregation that can lead to better representation of the global state.
@@ 5 Conclusions and Future Work
In this work, we propose a new methodology that develops a decentralized decision-making framework for inventory management. Our framework leverages the inherent graph structure and reduces policy overfitting by Gaussian noise injection in the value function. Our approach also offers several advantages that enhance the efficiency and effectiveness of inventory systems.
@@
Secondly, our methodology overcomes information sharing constraints at an online level but trains the control policies in a collaborative framework. This ensures effective coordination and communication is present within the different entities of the inventory management system. The communication between entities is further enhanced by leveraging the inherent graph structure of a supply chain. This also shifts the costs from online to offline, resulting in a more efficient stochastic optimal control policy. The methodology results in a closed-loop solution rather than an open-loop optimization problem which enhances fast, real-time decision-making capabilities.
Overall we show that the addition of noise, particularly in a system with a larger number of agents, reduces policy overfitting. We emphasize the importance of balancing effective exploration and exploitation by performing a sensitivity analysis on the noise intensity introduced in the value function. Our results show that too little noise leads to insufficient exploration whilst too much noise will increase the variance in the value function, disrupting the learning process. One should note that the amount of noise added will vary on the complexity of the environment and the number of agents present in the system.
@@ Acknowledgements
Appendix A Hyperparameter Values
@@
FC1 size 128
FC2 size 128
Appendix B Supply Chain Network Environment Parameters

### v2

@@ 4.2 Sensitivity to Gaussian Noise
The standard deviation values were changed from 0.0, 0.1, 0.2, 0.5, 1.0 and 2.0. To evaluate the impact of noise intensity, we compare the mean reward across 20 evaluation episodes, with the standard deviation shown as a shaded region in Figure 12a. This helps assess both policy performance and stability across different noise levels. The performance increases up to a standard deviation of 0.5. After which, when the noise intensity increases, the overall performance decreases. This occurs as when the standard deviation of the Gaussian distribution is small, there’s insufficient noise, leading to a lack of regularization. This can result in overfitting, limited exploration, and a tendency to settle into local optima. This is particularly detrimental in scenarios with large state spaces, where effective exploration becomes crucial for agents to learn optimal policies. Conversely, when the standard deviation is 1.0 and 2.0 the overall performance becomes worse as seen in Figure 13(a). This occurs when the noise intensity is too high, excessive noise can introduce significant randomness, increasing the variance in the value function which disrupts the learning process which may hinder convergence.
@@
(a) Execution curves after training. The mean episode reward is shown across 20 evaluation episodes, with shaded standard deviation.
(b) Entropy change during training for three different noise levels, illustrating the impact of noise intensity on policy exploration
@@
4.3 Uncertainty
In this section, we investigate how uncertainty affects the robustness of the proposed framework and its ability to adapt to the complexities of inventory management. Uncertainty, as defined by the International Bureau of Weights and Measures, represents the inherent doubt in measurement results, which has critical implications for decision-making and operational efficiency. In inventory management, uncertainty arises from sources such as demand variability, supply chain disruptions, and lead time fluctuations. These uncertainties can result in overstocking or stockouts, affecting profitability and operational reliability.
In our inventory management model, demand follows a Poisson distribution with a mean $\lambda_{\text{d}}$. The empirical demand distributions were compared against the theoretical Poisson Probability Mass Function (PMF) as shown in Figure 14. The results show that demand uncertainty grows with $\lambda_{\text{d}}$, with higher values leading to an increased probability of extreme demand spikes.
Figure 14: Comparison of empirical demand distributions with the theoretical Poisson PMF for different values of $\lambda_{\text{d}}$.
To assess the robustness of our learned RL policies, we trained the agents with our developed methodology (Section 3.4) under $\lambda_{\text{d}}=5$ and tested it under varying demand conditions $\lambda_{\text{d}}=3,4,5,6,7$.
(a) Profitability under varying demand levels
(b) Inventory levels and backlog across demand levels
Figure 15: Mean values with standard deviations for profitability, inventory levels, and backlog under different demand configurations.
Profitability is sensitive to demand fluctuations, as shown in Figure 15(a). At low demand levels, the RL agent tends to overstock, resulting in suboptimal profits. However, as demand increases, the variance in profit remains relatively stable, suggesting the policy maintains robustness despite fluctuations.
As expected, backlog increases with higher $\lambda_{\text{d}}$ values, reflecting the agent’s limited capacity to fulfill all demand within each time step. Conversely, on-hold inventory decreases with increasing demand, showing an inverse relationship with backlog, as shown in Figure 15(b).
As demand uncertainty grows, the standard deviation of inventory and backlog levels increases, indicating greater volatility in stock availability. Despite this, profit variance remains stable, suggesting that the RL policies are robust and maintain predictable financial performance. This suggests that the agent is effectively adapting its inventory and backlog management to maximize profits, despite increased demand variability.
@@ 4.4 Scalability
Moreover, Figure 17, compares the mean training time per iteration in seconds both in terms of methodologies and number of agents. As expected, as the number of agents increases, the training time per iteration increases due to the increased complexity of the learning process. The methodologies leveraging the graph structure also have a higher mean training time per iteration because they require complex message passing between agents in the graph. G-MAPPO, P-GCN-MAPPO, Reg-P-GCN-MAPPO utilize a GNN which requires training alongside the policies. Therefore, extracting and processing information from the graph structure itself adds computational overhead compared to simpler MARL approaches. As the number of agents increases, the training time per iteration increases at a faster rate for the methodologies using GNNs as the communication and information processing becomes increasingly complex as the number of agents and complexity of the structure increases. However, Figure 17 shows that methods with an integrated pooling mechanism have faster training times than those without one. This further emphasizes the scalability advantage of integrating a pooling mechanism compared to a standard GNN architecture in a MARL framework. The integration of the pooling mechanism reduces the dimensionality of the input to the central critic, reducing the computational overhead which leads to faster training times. Therefore, as opposed to naïvely concatenating information, MARL frameworks that employ smart information aggregation techniques can offer several advantages: reduced computational complexity, improved scalability and potential for improved performance by efficient aggregation that can lead to better representation of the global state.
@@ 5 Conclusions and Future Work
In this work, we propose a new methodology that develops a decentralized decision-making framework for inventory management. Our framework leverages the inherent graph structure and offers several advantages that enhance the efficiency and effectiveness of inventory systems.
@@
Secondly, our methodology overcomes information sharing constraints at an online level but trains the control policies in a collaborative framework. This ensures effective coordination and communication is present within the different entities of the inventory management system. The communication between entities is further enhanced by leveraging the inherent graph structure of a supply chain. This also shifts the costs from online to offline, resulting in a more efficient online stochastic optimal control policy. The methodology results in a closed-loop solution rather than an open-loop optimization problem which enhances fast, real-time decision-making capabilities.
@@ Acknowledgements
Appendix A Neural Network Architectures
The methodology consists of three main modules: multiple agent-specific actor networks, a centralized critic network, and a graph-based module utilizing a Graph Convolutional Network (GCN).
The actor (policy) networks are composed of three fully connected layers. The centralized critic follows a similar structure but outputs a scalar value without a Tanh activation.The graph module employs a Graph Convolutional Network (GCN) to capture spatial dependencies between agents. The architecture details are summarized in Table 2.
Parameter Actor (Policy) Critic GCN
Framework PyTorch (RLlib) PyTorch (RLlib) PyTorch (PyG)
Architecture MLP (3 layers) MLP (3 layers) GCN (3 layers)
Input State features State & Action Node features + adjacency
Hidden Layers 128-128-128 256-256-256 64-64-Output
Activation ReLU, tanh (output) ReLU ReLU
Output Action distribution (mean & std), Q-value estimate Processed node embeddings
Optimizer Adam Adam Adam
Learning Rate $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Weight Decay $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
Table 2: Neural Network Architectures for Actor, Critic, and GCN
The GCN processes agent connectivity using an adjacency matrix and updates node representations through message passing. The output embeddings are integrated into the centralized critic for value estimation and influence the policy network.
Appendix B Hyperparameter Values
@@
Appendix C Supply Chain Network Environment Parameters